OpenStudy (theloshua):

what is the average rate of change of the function y = x2 on the interval [2,5]?

OpenStudy (e.mccormick):

Know another name for average rate of change?

OpenStudy (theloshua):

no...

OpenStudy (e.mccormick):

Slope.

OpenStudy (e.mccormick):

Find the slope of the line with those given x values as the x pints and use the y=x^2 to find the y parts. (x, y=x^2) for the given [2,5] so x=2 and x=5.

OpenStudy (theloshua):

huhn? lol

OpenStudy (e.mccormick):

Average rate of change is the slope of the line defined by the end points of the sample. So for this sample range: 

OpenStudy (e.mccormick):

THe slope of this line: